Each visor requires a total of $4.00 in direct materials that includes an adjustable closure that the company purchases from a supplier at a cost of $1.50 each. Shadee wants to have 30 closures on hand on May 1, 20 closures on May 31, and 25 closures on June 30. Additionally, Shadee’s fixed manufacturing overhead is $1,000 per month, and variable manufacturing overhead is $1.25 per unit produced.         


Required:
1.
Determine Shadee's budgeted cost of closures purchased for May and June. (Round your answers to 2 decimal places.)

             


2. Determine Shadee's budget manufacturing overhead for May and June. (Do not round your intermediate calculations. Round your answers to 2 decimal places.)
Answer #1

1.

May June
Units sold 600.00 800.00
Add Ending Inventory 20 25
Less beginning inventory 30 20
Closures purchased 590 805
Purchases cost per closure $1.50 $1.50
Budgeted cost of closures purchased $885 $1207.50

2.

May June
Fixed manufacturing overhead $1,000 $1,000
Variable manufacturing overhead 1.25 x 575 = 718.75 1.25 x 810 = 1012.50
Budget manufacturing overhead $1,719 $2,013
Answer #2

1)


May
June
Units Sold
600
800
Add Ending Units
600+50=650
800+60=860
Less Starting Units
650-75=575
860-50=810
Add Closure End Units575+20=595810+25=835
Less Closure Start Units
595-30=565
835-20=815
Multiply Price of Closure
565x1.5=847.50
815x1.5=1222.50


answered by: TD